### v3

@@ Abstract
keywords:
port-Hamiltonian system, passivity, stability, system transformation, Kalman-Yacubovich-Popov inequality, Lyapunov inequality, even pencil, quadratic eigenvalue problem
pacs:
MSC Classification] 93A30, 93B17, 93B11
[…]
Acknowledgments
CB was supported by Einstein Foundation Berlin, through an Einstein Visiting Fellowship. VM gratefully acknowledges the support by Deutsche Forschungsgemeinschaft (DFG) as part of the collaborative research center SFB TRR 154 (grant no. 239904186), excellence cluster MATH+, Project AA4-12, and by BMBF (grant no. 05M22KTB) through EKSSE. HX was partially supported by Alexander von Humboldt Foundation.
@@ 1.1 Port-Hamiltonian systems
The synthesis of system models that describe complex physical phenomena often involves the coupling of independently developed subsystems originating within different disciplines. Systematic approaches to coupling such diversely generated subsystems prudently follows a system-theoretic network paradigm that focuses on the transfer of energy, mass, and other conserved quantities among the subsystems. When the subsystem models themselves arise from variational principles, then the aggregate system typically has structural features that reflects underlying conservation laws and very often it may be characterized as a port-Hamiltonian (pH) system, see BeaMV19 ; BeaMXZ18 ; GolSBM03 ; MasSB92 ; MehM19 ; MehS23 ; OrtSMM01 ; OrtSME02 ; Sch04 ; Sch06 ; SchM95 ; SchM02 ; SchM13 ; SchM23 for some major references and MehU23 for a detailed survey covering also the case of descriptor systems. Although pH systems may be formulated within a more general framework, we restrict ourselves to input-state-output pH systems, which have the form
@@
[…]
which has a natural interpretation as asserting that the increase in internal energy of the system, as measured by ${\mathcal{H}}$, cannot exceed the total work done on the system. ${\mathcal{H}}({\bf x})$ is a storage function associated with the supply rate, ${\bf y}(t)^{\top}{\bf u}(t)$. In the language of system theory, (3) constitutes the property that (1) is a passive system ByrIW91 .
One may verify with elementary manipulations that the inequality in (3) is an immediate consequence of the inequality in (2), and holds even when the coefficient matrices ${\bf J}$, ${\bf R}$, ${\bf F}$, ${\bf P}$, ${\bf S}$, and ${\bf N}$ depend on ${\bf x}$ or explicitly on time $t$ (see, MasSB92 ) or, indeed (with care taken to define suitable operator domains), when they represent linear operators acting on infinite dimensional spaces JacZ12 ; SchM02 . Notice that with a null input, ${\bf u}(t)=0$, the dissipation inequality asserts that ${\mathcal{H}}({\bf x})$ is non-increasing along any unforced system trajectory. Thus, ${\mathcal{H}}({\bf x})$ defines a Lyapunov function for the unforced system, so pH systems are implicitly Lyapunov stable HinP05 . Similarly, ${\mathcal{H}}({\bf x})$ is non-increasing along any system trajectory that produces a null output, ${\bf y}(t)=0$, so pH systems also have Lyapunov stable zero dynamics ByrI03 .
Port-Hamiltonian systems constitute a class of systems that is closed under power-conserving interconnection. This means that port-connected pH systems produce an aggregate system that must also be pH. This aggregate system will then be guaranteed to be both stable and passive. Modeling with pH systems, thus, represents physical properties in such a way as to facilitate automated modeling Kle13 while encoding physical properties explicitly into the structure of the equations. This framework also provides a compelling motivation to identify and preserve pH structure whenever it is present in order to produce high quality reduced order surrogate models, see BeaG11 ; GugPBS12 ; PolS10 .
Remark 1.
The pH system class has been extended (e.g., in BeaMXZ18 ; MehM19 ; MehU23 ), to include input-state-output pH descriptor (pHDAE) systems, which take on a similar form (at least in the case of autonomous systems):
${\bf E}\dot{\bf x}&=\left({\bf J}-{\bf R}\right){\bf e}({\bf x})+({\bf F}-{\bf P}){\bf u}(t),\\ {\bf y}(t)&=\ ({\bf F}+{\bf P})^{T}{\bf e}({\bf x})+({\bf S}+{\bf N}){\bf u}(t)\end{aligned}\quad\mbox{ where }\nabla_{\!{\bf x}}{\mathcal{H}}({\bf x})={\bf E}^{T}{\bf e}({\bf x}),$ (4)
@@ 1.2 Transforming general LTI systems to pH form
with ${\bf A}\in\mathbb{R}^{n\times n}$, ${\bf B}\in\mathbb{R}^{n\times m}$, ${\bf C}\in\mathbb{R}^{m\times n}$, and ${\bf D}\in\mathbb{R}^{m\times m}$. Following the previous discussion leading to (3), the system (5) is passive if there exists a continuously differentiable storage function ${\mathcal{H}}:\mathbb{R}^{n}\rightarrow[0,\infty)$ such that (3) holds for all admissible inputs ${\bf u}$ Wil72b .
@@
It is well-known BeaMV19 ; FauMPSW22 how one can use equivalence transformations to transform general LTI systems to port-Hamiltonian systems having the form
@@
Although ${\bf V}$ and ${\bf T}$ need only be invertible to be candidate transformations, the freedom we have in choosing a final pH realization allows for ${\bf V}$ to be assumed orthogonal and ${\bf T}$ to be constructed from compositions of well-conditioned triangular transformations with orthogonal transformations.
@@ 2 Passive systems and pH realizations
Since pH systems are structurally passive CheGH22 , our starting point for the construction of transformations to pH form is the following characterization of passivity introduced in Wil72b for minimal linear time invariant systems. The system (5) is minimal if it is both controllable and observable. The system (5) (and more specifically, the pair of matrices $({\bf A},{\bf B})$ with ${\bf A}\in\mathbb{R}^{n\times n}$, ${\bf B}\in\mathbb{R}^{n\times m}$) is controllable if $\mathop{\rm rank}\nolimits\left[\begin{array}[]{cc}s{\bf I}-{\bf A}&{\bf B}\end{array}\right]=n$ for all $s\in\mathbb{C}$. Similarly, the system (5) (and the pair $({\bf A},{\bf C})$ with ${\bf A}\in\mathbb{R}^{n\times n}$, ${\bf C}\in\mathbb{R}^{m\times n}$) is observable if $\mathop{\rm rank}\nolimits\left[\begin{array}[]{c}s{\bf I}-{\bf A}\\ {\bf C}\end{array}\right]=n$ for all $s\in\mathbb{C}$.
Theorem 3 (Wil72b ).
@@
Using the transformations in (8), this result has an immediate consequence for pH realizations, i.e. equivalent representations in pH form, see also CheGH22 .
Corollary 4.
@@
$-2\ \left[\begin{array}[]{cc}{\bf T}^{\top}&0\\ 0&{\bf I}\end{array}\right]\ \left[\begin{array}[]{lc}{\bf R}&{\bf P}\\ {\bf P}^{\top}&{\bf S}\end{array}\right]\ \left[\begin{array}[]{cc}{\bf T}&0\\ 0&{\bf I}\end{array}\right]\ \leq 0,$
@@
Remark 5.
@@
The construction described above for minimal passive systems is well known, see e.g. BeaMV19 . In the next subsection we discuss how to do this for non-minimal passive systems.
@@ 2.2 Construction for non-minimal systems
In the previous subsection we have discussed the existence of a positive definite solution ${\bf Q}$ of (9) under the assumption of minimality of the system. However, such solutions may exist even if the system is not minimal. A detailed analysis of the intricate relationship between passive systems, port-Hamiltonian descriptor systems and the solvability of the KYP inequality has recently been presented in CheGH22 and an explicit construction is presented in WeiWS94 . This analysis is particularly important when the system matrices arise from an interpolatory realization or a model reduction process where the resulting systems may be non-minimal or otherwise very close approximations of non-minimal systems. In these circumstances, the computation of a pH representation may be very sensitive to small perturbations arising from measurement or round-off errors.
Consider the following example from BeaMV19 .
Example 6.
@@
Note that when a system of the form (5) is generated by an interpolatory realization or other model reduction strategies, it may only be a close approximation to a passive system even when the original system is passive. In this case the inequality (9) might not be solvable, but a solution may exist for an adjacent system obtained from a small perturbation of the coefficients ${\bf A},{\bf B},{\bf C},{\bf D}$. How to obtain the “best" such perturbation is an important research topic, see AlaBKMM11 ; BruS13 ; Gri04 and references therein. We develop such perturbations when needed in the course of our construction, and so (referencing the pH representation), we are able to replace a nearly passive system with a nearby passive one.
Remark 7.
Since the matrix inequality (9) typically has an infinite number of solutions, an important question is how best to use the freedom in the choice of the solution of (9). One natural goal might be to minimize the distance to either instability or non-passivity or to maximize other robustness measures by pursuing the so-called analytic center of the solution set, see BanMNV20 ; MehV20 . Characterizing the solution set of (9) and its relationship to different robustness measures remains currently a (mostly) open problem, but see MehX24 for recent partial results.
@@
and via Lyapunov’s theorem LanT85 , this guarantees in turn that the unforced system $\dot{\bf x}={\bf A}{\bf x}$ is stable, i.e., ${\bf A}$ has all eigenvalues in the closed left half plane and those on the imaginary axis are semisimple; if the inequality (12) is strict, then the system is asymptotically stable, i.e., all eigenvalues of ${\bf A}$ are in the open left half plane. Passivity is encoded in the solvability of the full matrix inequality (9); strict passivity occurs when the inequality is strict, i.e., if the dissipation inequality (3) is strict, see KotA10 for a detailed analysis.
Remark 8.
In order to characterize the boundary of the solution sets of the LMIs (9) and (12) one needs to study the case when either the inequalities in (9) and (12) are not strict or when the resulting solutions are only semidefinite (or both). Extreme points of the solution set of (9) have recently been characterized in MehX24 .
@@ 3 Lyapunov and Riccati inequalities
The solutions of Lyapunov and Riccati inequalities as they arise in the characterizations (9) and (12) are typically addressed through semidefinite programming, see BoyEFB94 . For (9), an explicit characterization for all possible linear system realizations can be found in AlpL11 . In WeiWS94 a constructive method was proposed to deal with the case when ${\bf D}+{\bf D}^{\top}$ is singular. In contrast to AlpL11 , we assume that a general linear state-space system is given. We derive a constructive step-by-step procedure to check whether the linear system is passive, and if so, characterize all possible positive definite solutions of (9) for transforming (5) to a pH system (6).
@@
The stability of ${\bf A}$ is a necessary condition for (12) and (9) which require that ${\bf T}{\bf A}{\bf T}^{-1}+{\bf T}^{-\top}{\bf A}^{\top}{\bf T}^{\top}\leq 0$, or equivalently, that the Lyapunov inequality (12) has a positive definite solution ${\bf Q}={\bf T}^{\top}{\bf T}$. It is well known that the equality case in (12) always has a positive definite solution if ${\bf A}$ is stable, see LanT85 . In the following we recall, see e.g. BoyEFB94 , a characterization of the complete set of solutions of the inequality case.
If ${\bf A}$ is stable, but not asymptotically stable, then since all eigenvalues on the imaginary axis must be semi-simple, the real Jordan form of ${\bf A}$, see e.g. HorJ85 , guarantees the existance of a nonsingular matrix ${\bf M}\in\mathbb{R}^{n\times n}$ such that
@@
and consider the determination of the block ${\bf Q}_{1}$ separately from determination of the other blocks. Let $\mathcal{W}(n_{1})$ be the set of symmetric positive semidefinite matrices $\Theta_{1}\in\mathbb{R}^{n_{1}\times n_{1}}$ with the property that $\Theta_{1}{\bf x}\neq 0$ for any eigenvector ${\bf x}$ of ${\bf A}_{1}$. Then for any $\Theta_{1}\in\mathcal{W}(n_{1})$ we define ${\bf Q}_{1}$ to be the unique symmetric positive definite solution of the Lyapunov equation ${\bf A}_{1}^{\top}{\bf Q}_{1}+{\bf Q}_{1}{\bf A}_{1}=-\Theta_{1}$, see LanT85 . The other matrices $\hat{\bf Q}_{j}$, $j=2,\ldots,r$ are chosen of the form
@@
We have the following characterization of the solution set of (12), see (BoyEFB94, , Section 2.5.2).
Lemma 9.
@@
Remark 10.
Lemma 9 relies on the computation of the real Jordan form of ${\bf A}$, which generally is problematic in finite precision arithmetic. For the numerical computation of the solution it is better to use the real Schur form, see GolV96 .
@@
By Corollary 4 we can characterize (at least in the minimal case) the existence of a transformation to pH form via the existence of a symmetric positive definite matrix ${\bf Q}$ solving the KYP linear matrix inequality
@@
${\bf M}_{1}{\bf A}_{1}{\bf M}_{1}^{-1}=\mathop{\rm diag}\nolimits({\color[rgb]{0,0,0}\widetilde{\bf A}_{1}}+{\bf B}_{11}\hat{\bf S}^{-1}{\bf C}_{11},\alpha_{2}{\bf J}_{2},\ldots,\alpha_{r}{\bf J}_{r}).$
@@
In order to characterize the solution of the Riccati inequality (20), we first have to identify what happens if the system is not minimal. To numerically check minimality, we can use the orthogonal version of the Kalman decomposition, Kai80 ; Van79 , see also MehX00 .
Lemma 11.
@@
Proof.
The proof is straightforward and omitted.
By the previous lemma we can check controllability and observability but it is well-known PaiV81 that these properties can also be read off from Lagrangian invariant subspaces (if they exist) of certain Hamiltonian matrices associated with the equality case in (17), respectively (20), see e.g. LanR95 .
Lemma 13.
@@
The following lemma is also well-known, see e.g. LanR95 ; Meh91 .
Lemma 14.
@@
Remark 15.
By Lemma 13 we have seen that if the pair $({\bf A},{\bf C})$ is not observable then the matrix ${\bf W}_{2}$ in (44) may or may not be invertible. Hence, if ${\bf W}_{1}$ is invertible and ${\bf W}_{2}$ is not, then a symmetric solution to (45) still exits but is only positive semidefinite. This gives a characterization of the boundary of the solution set of the matrix inequality (16), see also CheGH22 for a detailed discussion of the existence of solution to the matrix inequality (16) if the controllability and observability conditions are violated.
Lemma 14 shows that under the conditions of controllability and observability the Riccati equation (45) has a solution ${\bf Q}>0$ whenever the Hamiltonian matrix ${\bf H}$ has a Lagrangian invariant subspace. The existence of such an invariant subspace depends only on the purely imaginary eigenvalues of ${\bf H}$, e.g., FreMX02 . When such an invariant subspace exists, then there are many such invariant subspaces. Note that the eigenvalues of ${\bf H}$ are symmetric with respect to the imaginary axis in the complex plane. If (44) holds, then the union of the eigenvalues of ${\bf Z}$ and $-{\bf Z}^{\top}$ form the spectrum of ${\bf H}$. One particular choice is that the spectrum of ${\bf Z}$ is in the closed left half complex plane, another choice is that it is in the closed right half complex plane. The two corresponding solutions of the Riccati equation (45) are the minimal solution ${\bf Q}_{-}$ and the maximal solution ${\bf Q}_{+}$ and all other solutions of the Riccati equation lie (in the Loewner ordering of symmetric matrices) between these extremal solutions.
Example 16.
[…]
@@
has a solution ${\bf Q}_{11}>0$. This is equivalent to the fact that the equality case in (47) has a positive definite solution, see BoyEFB94 ; LanR95 . Since the corresponding Hamiltonian matrix is
@@
The spectrum of $\widetilde{\bf H}_{11}$ is the union of the spectra of the submatrices ${\bf H}_{11}$, ${\bf A}_{22}$, and $-{\bf A}_{22}$. Since ${\bf A}_{22}$ is asymptotically stable, if ${\bf H}_{11}$ does not have purely imaginary eigenvalues, so is $\widetilde{\bf H}_{11}$. Hence, $\widetilde{\bf H}_{11}$ has a Lagrangian invariant subspace. Consider the Riccati equation
@@
If ${\bf H}_{11}$ as given in (48) has purely imaginary eigenvalues (potentially including 00) and we assume that it has a Lagrangian invariant subspace, then (47), written with equality, still has a solution $Q_{11}^{0}>0$ such that all eigenvalues of ${\bf A}_{11}-{\bf B}_{1}\hat{\bf S}^{-1}({\bf C}_{1}-{\bf B}_{1}^{\top}{\bf Q}_{11}^{0})$ are in the closed left half complex plane. Let $\widetilde{\bf Q}_{11}^{0}=\left[\begin{array}[]{cc}{\bf Q}_{11}^{0}&0\\ 0&0\end{array}\right]\geq 0$. Then $\widetilde{\Psi}_{11}(\widetilde{\bf Q}_{11}^{0})=0.$ Subtracting this from (51) yields the Riccati equation
@@
In a more general framework we can choose such a perturbation to turn the inequality in (9) into an equality by adding a positive semidefinite matrix to the left hand side to guarantee the existence of positive definite solutions. The set of all positive semidefinite perturbations of this kind will then help to characterize the solution set of (9). To do this in detail is beyond the scope of this paper and has recently been investigated in MehX24 .
@@
Proof.
The proof follows from the explicit construction.
Remark 19.

### v4

@@ Abstract
Keywords: port-Hamiltonian system, passivity, stability, system transformation, Kalman-Yacubovich-Popov inequality, Lyapunov inequality, even pencil, quadratic eigenvalue problem.
AMS subject classification. 93A30, 93B17, 93B11.
@@ 1.1 Port-Hamiltonian systems
The synthesis of system models that describe complex physical phenomena often involves the coupling of independently developed subsystems originating within different disciplines. Systematic approaches to coupling such diversely generated subsystems prudently follows a system-theoretic network paradigm that focuses on the transfer of energy, mass, and other conserved quantities among the subsystems. When the subsystem models themselves arise from variational principles, then the aggregate system typically has structural features that reflects underlying conservation laws and very often it may be characterized as a port-Hamiltonian (pH) system, see [6, 7, 24, 42, 47, 49, 52, 53, 56, 57, 58, 59, 60, 61] for some major references and [48] for a detailed survey covering also the case of descriptor systems. Although pH systems may be formulated within a more general framework, we restrict ourselves to input-state-output pH systems, which have the form
@@
${\mathcal{H}}({\bf x}(t_{1}))-{\mathcal{H}}({\bf x}(t_{0}))\leq\int_{t_{0}}^{t_{1}}{\bf y}(t)^{\top}{\bf u}(t)\ dt,$ (3)
which has a natural interpretation as asserting that the increase in internal energy of the system, as measured by $\mathcal{H}$}, cannot exceed the total work done on the system. ${\mathcal{H}}({\bf x})$ is a storage function associated with the supply rate, ${\bf y}(t)^{\top}{\bf u}(t)$. In the language of system theory, (3) constitutes the property that (1) is a passive system [17].
One may verify with elementary manipulations that the inequality in (3) is an immediate consequence of the inequality in (2), and holds even when the coefficient matrices $\bf J$}, $\bf R$}, $\bf F$}, $\bf P$}, $\bf S$}, and $\bf N$} depend on $\bf x$} or explicitly on time $t$ (see, [42]) or, indeed (with care taken to define suitable operator domains), when they represent linear operators acting on infinite dimensional spaces [33, 59]. Notice that with a null input, ${\bf u}(t)=0$, the dissipation inequality asserts that ${\mathcal{H}}({\bf x})$ is non-increasing along any unforced system trajectory. Thus, ${\mathcal{H}}({\bf x})$ defines a Lyapunov function for the unforced system, so pH systems are implicitly Lyapunov stable [31]. Similarly, ${\mathcal{H}}({\bf x})$ is non-increasing along any system trajectory that produces a null output, ${\bf y}(t)=0$, so pH systems also have Lyapunov stable zero dynamics [16].
Port-Hamiltonian systems constitute a class of systems that is closed under power-conserving interconnection. This means that port-connected pH systems produce an aggregate system that must also be pH. This aggregate system will then be guaranteed to be both stable and passive. Modeling with pH systems, thus, represents physical properties in such a way as to facilitate automated modeling [35] while encoding physical properties explicitly into the structure of the equations. This framework also provides a compelling motivation to identify and preserve pH structure whenever it is present in order to produce high quality reduced order surrogate models, see [5, 28, 55].
Remark 1
The pH system class has been extended (e.g., in [7, 47, 48]), to include input-state-output pH descriptor (pHDAE) systems, which take on a similar form (at least in the case of autonomous systems):
$\begin{aligned} {\bf E}\dot{\bf x}&=\left({\bf J}-{\bf R}\right){\bf e}({\bf x})+({\bf F}-{\bf P}){\bf u}(t),\\ {\bf y}(t)&=\ ({\bf F}+{\bf P})^{T}{\bf e}({\bf x})+({\bf S}+{\bf N}){\bf u}(t)\end{aligned}\quad\mbox{ where }\nabla_{\!{\bf x}}{\mathcal{H}}({\bf x})={\bf E}^{T}{\bf e}({\bf x}),$ (4)
@@ 1.2 Transforming general LTI systems to pH form
with ${\bf A}\in\mathbb{R}^{n\times n}$, ${\bf B}\in\mathbb{R}^{n\times m}$, ${\bf C}\in\mathbb{R}^{m\times n}$, and ${\bf D}\in\mathbb{R}^{m\times m}$. Following the previous discussion leading to (3), the system (5) is passive if there exists a continuously differentiable storage function ${\mathcal{H}}:\mathbb{R}^{n}\rightarrow[0,\infty)$ such that (3) holds for all admissible inputs $\bf u$} [64].
@@
It is well-known [6, 22] how one can use equivalence transformations to transform general LTI systems to port-Hamiltonian systems having the form
@@
Although $\bf V$} and $\bf T$} need only be invertible to be candidate transformations, the freedom we have in choosing a final pH realization allows for $\bf V$} to be assumed orthogonal and $\bf T$} to be constructed from compositions of orthogonal and well-conditioned triangular transformations.
@@ 2 Passive systems and pH realizations
Since pH systems are structurally passive [18], our starting point for the construction of transformations to pH form is the following characterization of passivity introduced in [64] for minimal linear time invariant systems. The system (5) is minimal if it is both controllable and observable. The system (5) (and more specifically, the pair of matrices $({\bf A},{\bf B})$ with ${\bf A}\in\mathbb{R}^{n\times n}$, ${\bf B}\in\mathbb{R}^{n\times m}$) is controllable if $\mathop{\rm rank}\nolimits\left[\begin{array}[]{cc}s{\bf I}-{\bf A}&{\bf B}\end{array}\right]=n$ for all $s\in\mathbb{C}$. Similarly, the system (5) (and the pair $({\bf A},{\bf C})$ with ${\bf A}\in\mathbb{R}^{n\times n}$, ${\bf C}\in\mathbb{R}^{m\times n}$) is observable if $\mathop{\rm rank}\nolimits\left[\begin{array}[]{c}s{\bf I}-{\bf A}\\ {\bf C}\end{array}\right]=n$ for all $s\in\mathbb{C}$.
Theorem 3 ([64])
@@
Using the transformations in (8), this result has an immediate consequence for pH realizations, i.e. equivalent representations in pH form, see also [18].
Corollary 4
@@
this gives (2) and ${\bf J},\,{\bf R},\,{\bf Q},\,{\bf F},\,{\bf P},\,{\bf S},\,\mbox{and }{\bf N}$ as defined in (10) will indeed determine a pH system.
Remark 5
@@
The construction described above for minimal passive systems is well known, see e.g. [6]. In the next subsection we discuss how to do this for non-minimal passive systems.
@@ 2.2 Construction for non-minimal systems
In the previous subsection we have discussed the existence of a positive definite solution $\bf Q$} of (9) under the assumption of minimality of the system. However, such solutions may exist even if the system is not minimal. A detailed analysis of the intricate relationship between passive systems, port-Hamiltonian descriptor systems and the solvability of the KYP inequality has recently been presented in [18] and an explicit construction is presented in [63]. This analysis is particularly important when the system matrices arise from an interpolatory realization or a model reduction process where the resulting systems may be non-minimal or otherwise very close approximations of non-minimal systems. In these circumstances, the computation of a pH representation may be very sensitive to small perturbations arising from measurement or round-off errors.
Consider the following example from [6].
Example 6
@@
Note that when a system of the form (5) is generated by an interpolatory realization or other model reduction strategies, it may only be a close approximation to a passive system even when the original system is passive. In this case the inequality (9) might not be solvable, but a solution may exist for an adjacent system obtained from a small perturbation of the coefficients ${\bf A},{\bf B},{\bf C},{\bf D}$. How to obtain the “best" such perturbation is an important research topic, see [1, 14, 27] and references therein. We develop such perturbations when needed in the course of our construction, and so (referencing the pH representation), we are able to replace a nearly passive system with a nearby passive one.
Remark 7
Since the matrix inequality (9) typically has an infinite number of solutions, an important question is how best to use the freedom in the choice of the solution of (9). One natural goal might be to minimize the distance to either instability or non-passivity or to maximize other robustness measures by pursuing the so-called analytic center of the solution set, see [4, 46]. Characterizing the solution set of (9) and its relationship to different robustness measures remains currently a (mostly) open problem, but see [51] for recent partial results.
@@
and via Lyapunov’s theorem [41], this guarantees in turn that the unforced system $\dot{\bf x}={\bf A}{\bf x}$ is stable, i.e., $\bf A$} has all eigenvalues in the closed left half plane and those on the imaginary axis are semisimple; if the inequality (12) is strict, then the system is asymptotically stable, i.e., all eigenvalues of $\bf A$} are in the open left half plane. Passivity is encoded in the solvability of the full matrix inequality (9); strict passivity occurs when the inequality is strict, i.e., if the dissipation inequality (3) is strict, see [36] for a detailed analysis.
Remark 8
In order to characterize the boundary of the solution sets of the LMIs (9) and (12) one needs to study the case when either the inequalities in (9) and (12) are not strict or when the resulting solutions are only semidefinite (or both). Extreme points of the solution set of (9) have recently been characterized in [51].
@@ 3 Lyapunov and Riccati inequalities
The solutions of Lyapunov and Riccati inequalities as they arise in the characterizations (9) and (12) are typically addressed through semidefinite programming, see [12]. For (9), an explicit characterization for all possible linear system realizations can be found in [2]. In [63] a constructive method was proposed to deal with the case when ${\bf D}+{\bf D}^{\top}$ is singular. In contrast to [2], we assume that a general linear state-space system is given. We derive a constructive step-by-step procedure to check whether the linear system is passive, and if so, characterize all possible positive definite solutions of (9) for transforming (5) to a pH system (6).
@@
The stability of $\bf A$} is a necessary condition for (12) and (9) which require that ${\bf T}{\bf A}{\bf T}^{-1}+{\bf T}^{-\top}{\bf A}^{\top}{\bf T}^{\top}\leq 0$, or equivalently, that the Lyapunov inequality (12) has a positive definite solution ${\bf Q}={\bf T}^{\top}{\bf T}$. It is well known that the equality case in (12) always has a positive definite solution if $\bf A$} is stable, see [41]. In the following we recall, see e.g. [12], a characterization of the complete set of solutions of the inequality case.
If $\bf A$} is stable, but not asymptotically stable, then since all eigenvalues on the imaginary axis must be semi-simple, the real Jordan form of $\bf A$}, see e.g. [32], guarantees the existance of a nonsingular matrix ${\bf M}\in\mathbb{R}^{n\times n}$ such that
@@
and consider the determination of the block ${\bf Q}_{1}$ separately from determination of the other blocks. Let $\mathcal{W}(n_{1})$ be the set of symmetric positive semidefinite matrices $\Theta_{1}\in\mathbb{R}^{n_{1}\times n_{1}}$ with the property that $\Theta_{1}{\bf x}\neq 0$ for any eigenvector $\bf x$} of ${\bf A}_{1}$. Then for any $\Theta_{1}\in\mathcal{W}(n_{1})$ we define ${\bf Q}_{1}$ to be the unique symmetric positive definite solution of the Lyapunov equation ${\bf A}_{1}^{\top}{\bf Q}_{1}+{\bf Q}_{1}{\bf A}_{1}=-\Theta_{1}$, see [41]. The other matrices $\hat{\bf Q}_{j}$, $j=2,\ldots,r$ are chosen of the form
@@
We have the following characterization of the solution set of (12), see [12, Section 2.5.2].
Lemma 9
@@
Remark 10
Lemma 9 relies on the computation of the real Jordan form of $\bf A$}, which generally is problematic in finite precision arithmetic. For the numerical computation of the solution it is better to use the real Schur form, see [25].
@@
Using Corollary 4 we can characterize (at least in the minimal case) the existence of a transformation to pH form via the existence of a symmetric positive definite matrix $\bf Q$} solving the KYP linear matrix inequality
@@
${\bf M}_{1}{\bf A}_{1}{\bf M}_{1}^{-1}=\mathop{\rm diag}\nolimits({\widetilde{\bf A}_{1}}+{\bf B}_{11}\hat{\bf S}^{-1}{\bf C}_{11},\alpha_{2}{\bf J}_{2},\ldots,\alpha_{r}{\bf J}_{r}).$
@@
In order to characterize the solution of the Riccati inequality (20), we first have to identify what happens if the system is not minimal. To numerically check minimality, we can use the orthogonal version of the Kalman decomposition, [34, 62], see also [50].
Lemma 11
@@
Proof. The proof is straightforward and omitted.
Using the previous lemma we may check controllability and observability, though it is well-known [54] that these properties can also be read off from Lagrangian invariant subspaces (if they exist) of certain Hamiltonian matrices associated with the equality case in (17), respectively (20), see e.g. [40].
Lemma 13
@@
The following lemma is also well-known, see e.g. [40, 44].
Lemma 14
@@
Remark 15
By Lemma 13 we have seen that if the pair $({\bf A},{\bf C})$ is not observable then the matrix ${\bf W}_{2}$ in (44) may or may not be invertible. Hence, if ${\bf W}_{1}$ is invertible and ${\bf W}_{2}$ is not, then a symmetric solution to (45) still exits but is only positive semidefinite. This gives a characterization of the boundary of the solution set of the matrix inequality (16), see also [18] for a detailed discussion of the existence of solution to the matrix inequality (16) if the controllability and observability conditions are violated.
Lemma 14 shows that under the conditions of controllability and observability the Riccati equation (45) has a solution ${\bf Q}>0$ whenever the Hamiltonian matrix $\bf H$} has a Lagrangian invariant subspace. The existence of such an invariant subspace depends only on the purely imaginary eigenvalues of $\bf H$}, e.g., [23]. When such an invariant subspace exists, then there are many such invariant subspaces. Note that the eigenvalues of $\bf H$} are symmetric with respect to the imaginary axis in the complex plane. If (44) holds, then the union of the eigenvalues of $\bf Z$} and $\bf Z$$^{\top}$ form the spectrum of $\bf H$}. One particular choice is that the spectrum of $\bf Z$} is in the closed left half complex plane, another choice is that it is in the closed right half complex plane. The two corresponding solutions of the Riccati equation (45) are the minimal solution $\bf Q$$_{-}$ and the maximal solution $\bf Q$$_{+}$ and all other solutions of the Riccati equation lie (in the Loewner ordering of symmetric matrices) between these extremal solutions.
Example 16
Consider the example ${\bf B}=\hat{\bf S}=1$, ${\bf C}=-1$, ${\bf A}=-1-\alpha$, where $\alpha>0$. So $\bf A$} and ${\bf A}-{\bf B}\hat{\bf S}^{-1}{\bf C}=-\alpha$ are both asymptotically stable. The Riccati equation (45) is $q^{2}-2\alpha q+1=0$, which does not have a real positive semidefinite solution when $\alpha\in(0,\,1)$. If $\alpha=1$, then it has a unique solution $q=1$ associated with ${\bf Z}=0$. If $\alpha>1$, then it has two solutions $\alpha\pm\sqrt{\alpha^{2}-1}>0$ with ${\bf Z}=\pm\sqrt{\alpha^{2}-1}$.
@@
has a solution ${\bf Q}_{11}>0$. This is equivalent to the fact that the equality case in (47) has a positive definite solution, see [12, 40]. Since the corresponding Hamiltonian matrix is
@@
The spectrum of $\widetilde{\bf H}_{11}$ is the union of the spectra of the submatrices ${\bf H}_{11}$, ${\bf A}_{22}$, and ${\bf A}_{22}$. Since ${\bf A}_{22}$ is asymptotically stable, if ${\bf H}_{11}$ does not have purely imaginary eigenvalues, so is $\widetilde{\bf H}_{11}$. Hence, $\widetilde{\bf H}_{11}$ has a Lagrangian invariant subspace. Consider the Riccati equation
@@
If ${\bf H}_{11}$ as given in (48) has purely imaginary eigenvalues (potentially including 0) and we assume that it has a Lagrangian invariant subspace, then (47), written with equality, still has a solution $Q_{11}^{0}>0$ such that all eigenvalues of ${\bf A}_{11}-{\bf B}_{1}\hat{\bf S}^{-1}({\bf C}_{1}-{\bf B}_{1}^{\top}{\bf Q}_{11}^{0})$ are in the closed left half complex plane. Let $\widetilde{\bf Q}_{11}^{0}=\left[\begin{array}[]{cc}{\bf Q}_{11}^{0}&0\\ 0&0\end{array}\right]\geq 0$. Then $\widetilde{\Psi}_{11}(\widetilde{\bf Q}_{11}^{0})=0.$ Subtracting this from (51) yields the Riccati equation
@@
In a more general framework we can choose such a perturbation to turn the inequality in (9) into an equality by adding a positive semidefinite matrix to the left hand side to guarantee the existence of positive definite solutions. The set of all positive semidefinite perturbations of this kind will then help to characterize the solution set of (9). To do this in detail is beyond the scope of this paper and has recently been investigated in [51].
@@
Proof. The proof follows from the explicit construction.
Remark 19